DBF borrows $9B by issuing 2-year bonds. ECB's cost of debt is 12%, so it will need to pay interest each year for the next 2 years, and then repay the principal $9B in year 2. ECB's marginal tax rate will remain 35% throughout this period. By how much does the interest tax shield increase the value of DBF?
Answer #1

Interest tax shield in Year 1 = Debt * Interest * Tax rate = $ 9 B * 12% * 35% = $ 0.378 B

Interest tax shield in Year 2 = $ 0.378 B

Increase in value of DBF = PV of interest tax shield = 0.378/(1+12%) + 0.378/(1+12%)^2 = $ 638,839.28 M